Problem 9-8A (Part Level Submission) In recent years, Wildhorse Co. has purchased three machines. Because of frequent employee turnover in the accounting department, a different accountant was in charge of selecting the depreciation method for each machine, and various methods have been used. Information concerning the machines is summarized in the table below. Salvage Value Useful Life (in years) 10 Depreciation Method Machine Acquired Jan. 1, 2015 July 1, 2016 Nov. 1, 2016 Cost $123,000 $23,000 11,400 6,700 Straight-line Declining-balance Units-of-activity 92,500 75,700 6 For the declining-balance method, Wildhorse Co. uses the double-declining rate. For the units-of-activity method, total machine hours are expected to be 34,500. Actual hours of use in the first 3 years were 2016, 850: 2017, 6,200: and 2018, 7,900 ▼ (a) Compute the amount of accumulated depreciation on each machine at December 31, 2018 MACHINE 1 MACHINE 2 MACHINE 3 Accumulated Depreciation at December 31 Click if you would like to Show Work for this question: Open Show Worlk
Answer #1

Answer

  • All working forms part of the answer
  • Workings

---Machine #1

A

Cost

$          123,000.00

B

Residual Value

$            23,000.00

C=A - B

Depreciable base

$          100,000.00

D

Life [in years]

10

E=C/D

Annual SLM depreciation each year

$            10,000.00

---Machine #2

A

Cost

$            92,500.00

B

Residual Value

$            11,400.00

C=A - B

Depreciable base

$            81,100.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$            16,220.00

F=E/C

SLM Rate

20.00%

G=F x 2

DDB Rate

40.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

2016 [6 months]

$              92,500.00

40.00%

$         18,500.00 [92500 x 40% x 6/12]

$             74,000.00

2017

$              74,000.00

40.00%

$         29,600.00

$             44,400.00

2018

$              44,400.00

40.00%

$         17,760.00

$             26,640.00

---Machine #3

A

Cost

$            75,700.00

B

Residual Value

$              6,700.00

C=A - B

Depreciable base

$            69,000.00

D

Usage

                     34,500 machine hours

E = C/D

Depreciation per machine hours

$                       2.00

  • Answers

Machine #1

Machine #2

Machine #3

Accumulated Depreciation at Dec 31 2018

$                10,000

$               17,760

$              15,800

Workings

[See Working for machine #1]

[See working for machine #2]

[7900 machine hours x $ 2 per machine hour]